Ray of light incident at an angle 60° on a √3 cm thick plate the shift in the path of the ray as it emerges out from the plate is? ( Refractive index of plate is √3)
1
Expert's answer
2019-07-05T09:13:39-0400

Initial conditions

"\alpha=60; d =\sqrt{3}; n=\sqrt{3}"

Find refraction angle

"\beta=arcsin(\frac{sin(\alpha)}{n})""\beta=30"

For shift we have

"\Delta x=\frac{dsin(\alpha-\beta)}{cos(\beta)}""\Delta x= 1 (cm)"